OpenStudy (anonymous):

Solve for x. x+1/2 - x+2/3 = x+3/4 a) x = 5 b) x = 7 c) x = -11 d) x = 23

OpenStudy (luigi0210):

Get common denominators

OpenStudy (anonymous):

12

OpenStudy (luigi0210):

Good now multiply everything by 12

OpenStudy (anonymous):

What exactly do you mean by everything

OpenStudy (anonymous):

x+12/2 x+24/3 x+36/4

OpenStudy (luigi0210):

\[12(x+\frac{1}{2}-x+\frac{2}{3})=12(x+\frac{3}{4})\]